OpenStudy (anonymous):

please help! find the vertex, focus,and directrix of the parabola and sketch its graph (x+3/2)^2=4(y-2)

OpenStudy (anonymous):

are they as the following vertex (-3/2,2) focus (-3/2,3) directrix is y-1 =0

OpenStudy (anonymous):

how Exactly did u get that??

OpenStudy (anonymous):

are the ans correct

OpenStudy (anonymous):

yeah according to the back of my book :)

OpenStudy (anonymous):

u hv to use shifting of the origin that is x+3/2=x' and y-2=y' and then use standard formulas for parabola

OpenStudy (anonymous):

c thats wat confuses me